Accueil > Sites heberges > CFHTLS-Deep > 2005 runs

 

2005 runs

Par Genevieve Soucail - 30/04/2007

 

Summary of 2005 runs

  • Summary of queue run 05BQ11 : 21 Dec. - 07 Jan.
    • A fabulous run, with only a few hours lost to bad seeing ! no bad conditions, no technical problems. A validation rate among the highest ever obtained with MegaCam. A lot of good data for CFHTLS.
    • D1 observed in 4 filters for a total of 7.3 hours
    • D2 observed in 4 filters for 9.2 hours.
    • D3 observed in 4 filters for 2.7 hours (beginning of observability)

  • Observations of D4 are closed (Dec. 15, 2005). See the D4 web page for details
  • Summary of queue run 05BQ09 : 23 Nov - 08 Dec.
    • A median run, with a bit more than one night lost to technical problems (electronics and shutter failure). About 4 nights lost to weather : 3 nights completely lost and marginal observing conditions for a total of one night spread over the run.
    • D4 observed in 4 filters for a total of 4.0 hours
    • D1 observed in 4 filters for 6.7 hours.
    • D2 observed in 4 filters for 6.4 hours.

  • Summary of queue run 05BQ07 : 22 Oct - 08 Nov.
    • A reasonable run, with 3 nights totally lost to weather, and about 3 more nights lost to clouds and spread over the full run.
    • D4 observed in 5 filters for a total of 9.2 hours.
    • D1 observed in 4 filters for 7.6 hours.
    • D2 started in 3 filters for 2 hours.

  • Summary of queue run 05BQ05 : 25 Sep - 11 Oct.
    • A good run, with 3 nights lost to weather and some cirrus thoughout a few nights spread over the run.
    • D4 observed in 5 filters for a total of 11.7 hours
    • D1 observed in 5 filters for 12.3 hours.

  • Summary of queue run 05BQ03 : 26 Aug - 11 Sep.
    • A good run, although one night was lost to technical problesm, and another one was done in half-mosaics mode (good data for Supernovae, but not validated for the Deep. Marginal observing conditions for a total of one night spread over the run
    • D4 observed in 4 filters for a total of 10.9 hours
    • D1 observed in 4 filters for 10.6 hours.

  • Observations of D3 are closed (Aug. 15, 2005). See the D3 web page for details

  • Summary of queue runs 05AQ07 & 05BQ01 : 24 Jul - 12 Aug.
    • Another average run. Very limited technical problems, and a new optical set-up with an increase of image quality. It seems stabilized for now. The beginning of the run had very bad seeing conditions, with very few validated exposures acquired during the 6 first nights for the Deep. Then seeing was OK, but a few more nights were lost to humidity.
    • D3 observed in g’ and r’ only in validated conditions (about 1h in total)
    • D4 observed in the 5 filters for a total of 9 hours
    • D1 observed in g’r’i’ for 2.7 hours.

  • Summary of queue run 05AQ06 : 29 Jun - 15 Jul.
    • An average run. Observations of CFHTLS perturbed by the Deep Impact event (occured on July 3rd) and the follow-ups of Temple1. Technical problems with the electronics of the camera and some bad weather (fog, cirrus and bad seeing) led to 25% of the time lost.
    • Autofocus is working and only a very few focus sequences are taken each night.
    • D4 observed in the 5 filters for a total of 10 hours
    • D3 observed in g’r’i’z’ for 8.8 hours.

  • Summary of queue run 05AQ05 : 27 May - 16 Jun.
    • A good run. No technical problems. Some time lost to weather, cirrus or bad IQ, within the average range.
    • The last night of the run completed without any focus sequence. Autofocus is working !
    • D3 observed in the 5 filters for a total of 10.5 hours
    • Beginning of D4 observations. 2.3 hours in r’ and i’ only.

  • Observations of D2 are closed (May. 20, 2005). See the D2 web page for details

  • Summary of queue run 05AQ04 : 04 May - 18 May
    • The best QSO run of the semester ! No time lost to weather, photometric sky most of the time and IQ below the specs nearly all the time. About 1.5 night lost to some technical problems on the telescope.
    • D2 observed in 2 filters (r’ and i’). End of its observability
    • D3 observed in good conditions in 4 filters for a total of 11.2 hours !

  • Summary of queue run 05AQ03 : 31 March - 16 April
    • A significant improvement of weather side compared to preivous runs, although the first half of the run was limited due to non-photometric or high IQ periods. A few technical problems with the camera readout leading to a few hours lost.
    • D2 observed for 5.1 hours (validated) and D3 for 6.3 hours.
    • Both fields observed in 4 filters only, under the recommendation of the Steering Group.

  • Summary of queue run 05AQ02 : 10 March - 20 March
    • A short run (11 nights only), again plagued by bad weather. Only 4 nights available for observations
    • D2 observed for 2.9 hours (validated) and D3 for 3.8 hours.
    • D2 observed in 4 filters, D3 in 5 (u* observations obtained in marginal seeing conditions, but validated).

  • Summary of queue run 05AQ01 : 01 February - 17 February
    • Another a bad run with 11 nights lost to weather (ice, wind) or with seeing out of range for LS.
    • D2 observed for 4.3 hours (validated) and D3 for 2.2 hours.
    • Only g’, r’, i’, z’ data acquired, a few in good seeing conditions.

back to recent updates

 

 

[Page Précédente] [Dans la même rubrique] [Sommaire]

 

 

Site du CNRS Logo OMP
Logo UPS
Logo INSU-CNRS

Annuaire

Rechercher

Sur ce site



Bibliotheque

Laboratoire Astrophysique de Toulouse - Tarbes (UMR5572)

CNRS (Midi-Pyrénées)

Univ. Paul Sabatier

Liens utiles

Accueil Imprimer Contact mail Plan du site Crédits